# DeepPolar: Inventing Nonlinear Large-Kernel Polar Codes via Deep Learning

### v1

@@ Abstract
Polar codes, developed on the foundation of Arikan’s polarization kernel, represent a breakthrough in coding theory and have emerged as the state-of-the-art error-correction-code in short-to-medium block length regimes. Importantly, recent research has indicated that the reliability of polar codes can be further enhanced by substituting Arikan’s kernel with a larger one, leading to a faster polarization. However, for short-to-medium block length regimes, the development of polar codes that effectively employ large kernel sizes has not yet been realized. In this paper, we explore a novel, non-linear generalization of polar codes with an expanded kernel size, which we call DeepPolar codes. Our results show that DeepPolar codes effectively utilize the benefits of larger kernel size, resulting in enhanced reliability compared to both the existing neural codes and conventional polar codes. Source code is available at this link.
@@ 1 Introduction
Polar codes and SC decoding are optimal for large asymptotic blocklengths; however, practical finite-length performance is lacking. Recognizing this limitation, recent research has focused on augmenting the polar encoder and improving the decoding performance. Specifically, the concatenation of cyclic redundancy check (CRC) with polar codes improves distance properties. Combining this with successive cancellation list (SCL) decoding markedly enhances decoding performance (Niu & Chen, 2012). Consequently, Polar codes have emerged as the state-of-the-art in the short-to-medium block length regime, leading to their inclusion in 5G standards. Nonetheless, SCL decoding introduces considerable decoding complexity and latency, marking a trade-off between performance and computational efficiency.
@@
An alternative idea involves increasing the size of the polarization kernel. Indeed, (Korada et al., 2010) prove that polarization holds for all kernels provided they are not unitary and not upper triangular under any column permutation. Further, one can find large polarization kernels ($\ell\geq 8$) that achieve faster polarization (characterized by the "scaling exponent"). Despite these theoretical advantages, binary Polar codes with large kernels exhibit poor performance at practical short-to-medium block lengths and face exponential increases in computational complexity with kernel size $\ell$.
Our research delves into the innovative intersection between algebraic coding theory and machine learning by exploring non-linear generalizations of polarization-driven structures. This can be achieved by parameterizing each kernel by a neural network, combining the information-theoretic properties of polarization-driven code structures with the adaptability and learning capabilities of deep learning. This interplay between algebraic coding structures and deep learning is a relatively unchartered territory. Our algorithm builds upon the groundwork laid by (Makkuva et al., 2021), which introduced KO codes as a non-linear generalization of RM codes. Although this work marked a significant leap forward, it is limited by its dependence on Reed-Muller (RM) encoding-decoding schemes, which restricted its applicability across a broader range of rates. Importantly, the architecture of KO codes proved inapplicable for Polar codes, being only scalable upto a (64,7) code. In contrast, our work proposes DeepPolar codes, a non-linear generalization of the encoding and decoding structures of Polar codes (which encompass RM codes as a special case). This allows us to scale seamlessly to various rates and block lengths.
DeepPolar(256,37)KO(8,2)Polar(256,37)RM(8,2)Kernel size $\ell$
@@ 2.1 Channel coding
Channel coding is a technique to add redundancy to the transmission to make it robust against noise added by the communication channel. More precisely, let $\boldsymbol{u}=(u_{0},\ldots,u_{k-1})\in\{0,1\}^{k}$ denote a block of information/message bits that we wish to transmit. A code consists of an encoder and decoder pair. The encoder $g_{\phi}:\{0,1\}^{k}\to\{0,1\}^{n}$ maps these message bits into a binary codeword $\boldsymbol{x}$ of length $n$, i.e. $\boldsymbol{x}=g(\boldsymbol{u})$. The codewords are mapped to real/complex values by modulation (eg, Binary Phase Shift Keying (BPSK)). The channel, denoted as $W_{Y|X}(\cdot|\cdot)$, corrupts the codeword $\boldsymbol{x}$ to its noisy version $\boldsymbol{y}\in\mathbb{R}^{n}$. Upon receiving the corrupted codeword, the decoder $f_{\theta}$ estimates the message bits as $\hat{\boldsymbol{u}}=f_{\theta}(\boldsymbol{y})$. The performance of the code is measured using standard error metrics such as Bit-Error-Rate (BER) or Block-Error-Rate (BLER): $\mathrm{BER}(g_{\phi},f_{\theta})\triangleq(1/k)\sum_{i}\mathbb{P}[\hat{\boldsymbol{u}}_{i}\neq\boldsymbol{u}_{i}]$, whereas $\mathrm{BLER}(g_{\phi},f_{\theta})\triangleq\mathbb{P}[\hat{\boldsymbol{u}}\neq\boldsymbol{u}]$.
@@ 2.2 Polar codes
[…]
@@
Utilizing this structure, the encoding can be efficiently performed via a recursive coordinate-wise application of the Plotkin transform on a binary tree, called the Plotkin tree. To encode a block of message bits $\boldsymbol{u}=(u_{0},\ldots,u_{k-1})\in\{0,1\}^{k}$, we first embed them into a source message vector $\boldsymbol{m}\triangleq(m_{1},\ldots,m_{n})=(0,\ldots,u_{0},0,\ldots,u_{1},0,\ldots,u_{k-1},0,\ldots)\in\{0,1\}^{n}$, where $\boldsymbol{m}_{I_{k}}=\boldsymbol{u}$ and $\boldsymbol{m}_{\mathcal{I}^{C}}=0$ for some $\mathcal{I}\subseteq[n]$. Since the message block $\boldsymbol{m}$ contains the information bits $\boldsymbol{u}$ only at the indices pertaining to $I_{k}$, the set $I_{k}$ is called the information set, and its complement $\mathcal{F}=\mathcal{I}^{C}$ the frozen set. We describe the encoding on a Plotkin tree via a small example, Polar(4,3) - illustrated in  Fig. 3(a). Here, $\mathcal{F}$={0}. Consider an input of size $k=3$, $u=[u_{0},u_{1},u_{2}]$. At the input level (depth 1), we freeze $m_{0}$, i.e., $m_{0}=0$ and assign $\boldsymbol{u}$ to the remaining positions. Applying the Plotkin transform, $(0,u_{0})\to(0\oplus u_{0},u_{0})$ and $(u_{1},u_{2})\to(u_{1}\oplus u_{2},u_{2})$. At the second level, we apply the same operation coordinatewise to these vectors i.e, $((0\oplus u_{0}),(u_{1}\oplus u_{2}))\to((0\oplus u_{0})\oplus(u_{1}\oplus u_{2}),u_{1}\oplus u_{2})$ and $(u_{0},u_{2})\to(u_{0}\oplus u_{2},u_{2})$. The final encoded vector is the concatenation of the outputs from the second-level nodes i.e, $(u_{0}\oplus u_{1}\oplus u_{2},u_{0}\oplus u_{2},u_{1}\oplus u_{2},u_{2})$. For a general $(n,k)$ polar code, the encoding proceeds similarly up to $m=\log_{2}{n}$ levels.
@@
$L_{i}=log\left(\frac{\mathbb{P}(u_{i}=0|(y,\hat{u}^{(i-1)})}{\mathbb{P}(u_{i}=1|(y,\hat{u}^{(i-1)})}\right).$ (1)
@@
Large kernel Polar codes. Polar codes are optimal (capacity-achieving) at asymptotic blocklengths due to the phenomenon of channel polarization. (Korada et al., 2010) prove that replacing the conventional $2\times 2$ kernel with an $\ell\times\ell$ binary kernel ($\ell>2$) still results in polarization, provided this matrix is non-singular and not upper triangular under any column permutation. Further, large kernel polar codes that achieve capacity at shorter block lengths compared to conventional polar codes have been found, as indicated by better scaling exponents (Fazeli & Vardy, 2014; Fazeli et al., 2020). Notably, the kernel size must be expanded to 8 to surpass Arikan’s $G^{\otimes 2}$ kernel, as no size 4 linear kernel offers an improved scaling exponent. While this generally translates to better finite-length performance, the complexity of decoding grows exponentially with kernel size (Trofimiuk & Trifonov, 2021); and decoders with comparable computational complexity (eg, SC decoding) are highly sub-optimal at short-to-medium block lengths (App. B).
@@
In this work, we design a non-linear generalization of the polar coding and decoding structure by both expanding the kernel size and parameterizing them as neural networks. By expanding the design space to include non-linearity and larger kernels, we aim to discover more reliable codes within the Plotkin code family.
@@ 3.1 Proposed Architecture
(a) Stage 1 - Curriculum learning for kernels
(b) Stage 2 - Initialize encoder and decoder by pretrained kernels
@@
DeepPolar encoder. Building on the foundational encoding structure of Polar codes (Fig. 3a), which employs the recursive application of a binary kernel over a Plotkin tree, we introduce DeepPolar codes as a non-linear extension of this concept. The first major modification in our approach is expanding the conventional $2\times 2$ kernel to a larger $\ell\times\ell$ kernel. While this generalization does not inherently lead to superior Polar codes at short blocklengths compared to the standard Arikan kernel, as discussed in the previous section (Fig. 3b), it lays the groundwork for further innovations in our design.
On the other hand, we generalize the Plotkin transform by replacing each XOR operation with a 2-input NN $g_{d,b}$, where $d$ is the depth in the encoding tree, and $b$ is the index of the NN at depth $d$. A DeepPolar $(n,k,\ell)$ inherits the same parameters as a Polar$(n,k)$ code. The encoding structure is analogous to conventional polar codes and is illustrated in Fig. 3(c) - DeepPolar (4,3) with $\ell=2$. We use the same frozen set $\mathcal{F}$=0 as the conventional polar code. Applying the neural Plotkin transform, $(0,u_{0})\to v_{10}=(g_{01}(0,u_{0}),u_{0})$ and $(u_{1},u_{2})\to v_{11}=(g_{11}(u_{1},u_{2}),u_{2})$. At the second level, we apply the kernel $g_{20}$ coordinatewise to these vectors. This coordinatewise operation is the key inductive bias in DeepPolar encoding. The final encoded vector is the concatenation of the outputs from the second-level nodes i.e, $(g_{20}(v_{10},v_{11}),v_{11})$. For a general DeepPolar $(n,k,\ell=2)$code, the encoding proceeds in a similar way upto $m=\log_{2}{n}$ levels.
Our work synergizes these two directions by employing a non-linear kernel of size $\ell\times\ell$. Fig. 3(d) illustrates DeepPolar $(16,8,\ell=4)$, which mirrors the Plotkin-4 encoder for Polar$(16,8)$. The encoding proceeds similarly, with neural networks $g_{db}$ replacing the Plotkin mapping. A power constraint, $\|x\|^{2}=n$ is enforced at the output of the encoder. App. E.5 discusses the DeepPolar encoder architecture in detail.
While large kernels in the context of Polar codes fail to improve over the standard 2x2 kernel, a critical distinguishing factor of DeepPolar is that we parameterize the kernel at each internal node $i$ of the Plotkin tree by a neural network $g_{i}$. We hypothesize that this innovation unlocks performance gains through the expanded function space afforded by both non-linearization and the increased kernel size. Additionally, this formulation allows us to scale neural Polar codes to larger blocklengths and rates.
@@
DeepPolarKO(8,2)Polar(256,37)RM(8,2)DeepPolar-binarySignal-to-noise ratio (SNR) [dB]
@@
Curriculum. We address this by introducing a principled two-stage curriculum training procedure that capitalizes on the inherent nested hierarchy of Polar codes. We leverage two key properties of the Polar coding structure: (1)Hierarchy in k: A Polar$(n,k)$ subsumes all codewords of a $\text{Polar}(n,k)$ code subsumes all the codewords of lower-rate subcodes $\text{Polar}(n,i),1\leq i\leq k$. Curriculum training leveraging the hierarchy in $k$ improved both convergence speed and reliability in the context of neural polar decoding (Hebbar et al., 2023). In the first stage, we use a similar C2N curriculum to pretrain the parameters for each kernel. We progressively train encoder-decoder pairs for DeepPolar ($\ell,j,\ell)$ codes, for $j=1\to k$ as highlighted in Fig. 4a), denoted by $(g_{\phi},f_{\theta})(\ell,j)$. (2)Hierarchy in n: In the Plotkin tree, an $\ell\times\ell$ kernel is applied coordinatewise at each level of the Plotkin tree. In the second stage of our curriculum, we initialize each kernel in DeepPolar ($n$,$k$) encoder and decoder with the pre-trained kernels from the first stage, applied at all depths, as illustrated in Fig. 4b. For instance, $g_{12}$ and $g_{20}$, which have three non-frozen inputs, are initialized by a pre-trained kernel $g_{\phi}(4,3)$. Such a systematic and phased training approach significantly enhances the network’s learning efficiency and generalizability (Sec. 4.4).

### v2

@@ Abstract
Progress in designing channel codes has been driven by human ingenuity and, fittingly, has been sporadic. Polar codes, developed on the foundation of Arikan’s polarization kernel, represent the latest breakthrough in coding theory and have emerged as the state-of-the-art error-correction code for short-to-medium block length regimes. In an effort to automate the invention of good channel codes, especially in this regime, we explore a novel, non-linear generalization of Polar codes, which we call DeepPolar codes. DeepPolar codes extend the conventional Polar coding framework by utilizing a larger kernel size and parameterizing these kernels and matched decoders through neural networks. Our results demonstrate that these data-driven codes effectively leverage the benefits of a larger kernel size, resulting in enhanced reliability when compared to both existing neural codes and conventional Polar codes. Source code is available at this link.
@@ 1 Introduction
Polar codes and SC decoding are optimal for large asymptotic blocklengths; however, practical finite-length performance is lacking. Recognizing this limitation, recent research has focused on augmenting the polar encoder and improving the decoding performance. Specifically, the concatenation of cyclic redundancy check (CRC) with polar codes improves distance properties. Combining this with successive cancellation list (SCL) decoding markedly enhances decoding performance (Tal & Vardy, 2015; Niu & Chen, 2012). Consequently, Polar codes have emerged as the state-of-the-art in the short-to-medium block length regime, leading to their inclusion in 5G standards. Nonetheless, SCL decoding introduces considerable decoding complexity and latency, marking a trade-off between performance and computational efficiency.
@@
An alternative idea involves increasing the size of the polarization kernel. Indeed, (Korada et al., 2010) prove that polarization holds for all kernels provided they are not unitary and not upper triangular under any column permutation. Further, one can find large polarization kernels ($\ell\geq 8$) that achieve faster polarization (characterized by the "scaling exponent"). Despite these theoretical advantages, Polar codes with large kernels are not preferred in practice due to their increased decoding complexity. Nevertheless, there is recent work addressing these challenges (Trifonov, 2023) based on specially designed polarization kernels.
In this work, we aim to address the question: How can we automate the search for good codes? Indeed, it is possible by parameterizing and learning both the encoder and decoder using neural networks. However, constructing effective non-linear codes using this approach is highly challenging: it is well-documented in literature that naively parameterizing with off-the-shelf neural architectures often results in performance worse than even repetition codes, as elaborated further in App. F.3. Rather, a more promising approach is to design neural architectures that enable structured redundancy. Specifically, our work delves into the innovative intersection between algebraic coding theory and machine learning by exploring non-linear generalizations of polarization-driven structures. This can be achieved by parameterizing each kernel by a neural network, combining the information-theoretic properties of polarization-driven code structures with the adaptability and learning capabilities of deep learning. This interplay between algebraic coding structures and deep learning is a relatively unchartered territory. Our algorithm builds upon the groundwork laid by (Makkuva et al., 2021), who introduced KO codes as a non-linear generalization of RM codes. Although this work marked a significant leap forward, it is limited by its dependence on Reed-Muller (RM) encoding-decoding schemes, which restricted its applicability across a broader range of rates. Importantly, the architecture of KO codes proved inapplicable for Polar codes, being only scalable upto a (64,7) code. In contrast, our work proposes DeepPolar codes, a non-linear generalization of the encoding and decoding structures of Polar codes (which encompass RM codes as a special case). This allows us to scale seamlessly to various rates and block lengths.
DeepPolar(256,37)KO(8,2)Polar(256,37)RM(8,2Kernel size $\ell$
@@ 2.1 Channel coding
Channel coding is a technique to add redundancy to the transmission to make it robust against noise added by the communication channel. More precisely, let $\boldsymbol{u}=(u_{0},\ldots,u_{k-1})\in\{0,1\}^{k}$ denote a block of information/message bits that we wish to transmit. A code consists of an encoder and decoder pair. The encoder $g_{\phi}:\{0,1\}^{k}\to\{0,1\}^{n}$ maps these message bits into a binary codeword $\boldsymbol{x}$ of length $n$, i.e. $\boldsymbol{x}=g(\boldsymbol{u})$. The codewords are mapped to real/complex values by modulation (eg, Binary Phase Shift Keying (BPSK)). The channel, denoted as $W_{Y|X}(\cdot|\cdot)$, corrupts the codeword $\boldsymbol{x}$ to its noisy version $\boldsymbol{y}\in\mathbb{R}^{n}$. Upon receiving the corrupted codeword, the decoder $f_{\theta}$ estimates the message bits as $\hat{\boldsymbol{u}}=f_{\theta}(\boldsymbol{y})$. The performance of the code is measured using standard error metrics such as Bit-Error-Rate (BER) or Block-Error-Rate (BLER): $\mathrm{BER}(g_{\phi},f_{\theta})\triangleq(1/k)\sum_{i}\mathbb{P}[\hat{u}_{i}\neq u_{i}]$, whereas $\mathrm{BLER}(g_{\phi},f_{\theta})\triangleq\mathbb{P}[\hat{\boldsymbol{u}}\neq\boldsymbol{u}]$.
@@ 2.2 Polar codes
Figure 3: (a) Polar$(4,3)$ encoding structure using the standard $2\times 2$ kernel. Encoding is performed recursively on the Plotkin tree. (b) Polar$(16,8)$ encoding using $4\times 4$ kernels. (c) DeepPolar ($4,3,\ell=2)$ replaces the xor operation in Plotkin - $2\times 2$ by neural networks. (d) DeepPolar ($16,8,\ell=4)$ : Scaling the DeepPolar encoding to higher-order kernels enables us to achieve good reliability. We are the first to explore this design space
@@
Utilizing this structure, the encoding can be efficiently performed via a recursive coordinate-wise application of the Plotkin transform on a binary tree, called the Plotkin tree. To encode a block of message bits $\boldsymbol{u}=(u_{0},\ldots,u_{k-1})\in\{0,1\}^{k}$, we first embed them into a source message vector $\boldsymbol{m}\triangleq(m_{1},\ldots,m_{n})=(0,\ldots,u_{0},0,\ldots,u_{1},0,\ldots,u_{k-1},0,\ldots)\in\{0,1\}^{n}$, where $\boldsymbol{m}_{\mathcal{I}}=\boldsymbol{u}$ and $\boldsymbol{m}_{\mathcal{I}^{C}}=0$ for some $\mathcal{I}\subseteq[n]$. Since the message block $\boldsymbol{m}$ contains the information bits $\boldsymbol{u}$ only at the indices pertaining to $\mathcal{I}$, the set $\mathcal{I}$ is called the information set, and its complement $\mathcal{F}=\mathcal{I}^{C}$ the frozen set. We describe the encoding on a Plotkin tree via a small example, Polar(4,3) - illustrated in  Fig. 3(a). Here, $\mathcal{F}$={0}. Consider an input of size $k=3$, $u=[u_{0},u_{1},u_{2}]$. At the input level (depth 1), we freeze $m_{0}$, i.e., $m_{0}=0$ and assign $\boldsymbol{u}$ to the remaining positions. Applying the Plotkin transform, $(0,u_{0})\to(0\oplus u_{0},u_{0})$ and $(u_{1},u_{2})\to(u_{1}\oplus u_{2},u_{2})$. At the second level, we apply the same operation coordinatewise to these vectors i.e, $((0\oplus u_{0}),(u_{1}\oplus u_{2}))\to((0\oplus u_{0})\oplus(u_{1}\oplus u_{2}),u_{1}\oplus u_{2})$ and $(u_{0},u_{2})\to(u_{0}\oplus u_{2},u_{2})$. The final encoded vector is the concatenation of the outputs from the second-level nodes i.e, $(u_{0}\oplus u_{1}\oplus u_{2},u_{0}\oplus u_{2},u_{1}\oplus u_{2},u_{2})$. For a general $(n,k)$ polar code, the encoding proceeds similarly up to $m=\log_{2}{n}$ levels.
@@
$L_{i}=\log\left(\frac{\mathbb{P}(u_{i}=0|y,\hat{\boldsymbol{u}}^{(i-1)})}{\mathbb{P}(u_{i}=1|(y,\hat{\boldsymbol{u}}^{(i-1)})}\right).$ (1)
@@
Large kernel Polar codes. Polar codes are optimal (capacity-achieving) at asymptotic blocklengths due to the phenomenon of channel polarization. (Korada et al., 2010) prove that replacing the conventional $2\times 2$ kernel with an $\ell\times\ell$ binary kernel ($\ell>2$) still results in polarization, provided this matrix is non-singular and not upper triangular under any column permutation. Further, large kernel polar codes that achieve capacity at shorter block lengths compared to conventional polar codes have been found, as indicated by better scaling exponents (Fazeli & Vardy, 2014; Fazeli et al., 2020). Notably, the kernel size must be expanded to 8 to surpass Arikan’s $G^{\otimes 2}$ kernel, as no size 4 linear kernel offers an improved scaling exponent. These kernels enable better finite-length performance at the cost of increased decoding complexity.
@@
In this work, we design a non-linear generalization of the polar coding and decoding structure by both expanding the kernel size and parameterizing them as neural networks. By expanding the design space to include non-linearity and larger kernels, we aim to discover more reliable codes within the neural Plotkin code family.
@@ 3.1 Proposed Architecture
(a)) Stage 1 - Curriculum learning for kernels
(b)) Stage 2 - Initialize encoder and decoder by pretrained kernels
@@
DeepPolar encoder. Building on the foundational encoding structure of Polar codes (Fig. 3a), which employs the recursive application of a binary kernel over a Plotkin tree, we introduce DeepPolar codes as a non-linear extension of this concept. The first major modification in our approach is expanding the conventional $2\times 2$ kernel to a larger $\ell\times\ell$ kernel. This is inspired by the existence of larger binary kernels that outperform the Arikan kernel (App. B.2.)
On the other hand, we generalize the Plotkin transform by replacing each XOR operation with a 2-input NN $g_{d,b}$, where $d$ is the depth in the encoding tree, and $b$ is the index of the NN at depth $d$. A DeepPolar $(n,k,\ell)$ inherits the same parameters as a Polar$(n,k)$ code. The encoding structure is analogous to conventional polar codes and is illustrated in Fig. 3(c) - DeepPolar (4,3) with $\ell=2$. We use the same frozen set =$\mathcal{F}=\{0\}$ as the conventional polar code. Applying the neural Plotkin transform, $(0,u_{0})\to v_{10}=(g_{10}(0,u_{0}),u_{0})$ and $(u_{1},u_{2})\to v_{11}=(g_{11}(u_{1},u_{2}),u_{2})$. At the second level, we apply the kernel $g_{20}$ coordinatewise to these vectors. This coordinatewise operation is the key inductive bias in DeepPolar encoding. The final encoded vector is the concatenation of the outputs from the second-level nodes i.e, $(g_{20}(v_{10},v_{11}),v_{11})$. For a general DeepPolar $(n,k,\ell=2)$code, the encoding proceeds in a similar way upto $m=\log_{2}{n}$ levels.
Our work synergizes these two directions by employing a non-linear kernel of size $\ell\times\ell$. Fig. 3(d) illustrates DeepPolar $(16,8,\ell=4)$, which mirrors the Plotkin-4 encoder for Polar$(16,8)$. The encoding proceeds similarly, with neural networks $g_{d,b}$ replacing the Plotkin mapping. A power constraint, $\|x\|^{2}=n$ is enforced at the output of the encoder. App. E.5 discusses the DeepPolar encoder architecture in detail.
DeepPolar leverages the Polar encoding structure by parameterizing the kernel at each internal node at depth $d$ and bit position $b$ of the Plotkin tree by a neural network $g_{d,b}$. We hypothesize that this innovation unlocks performance gains through the expanded function space afforded by both non-linearization and the increased kernel size. Additionally, this formulation allows us to scale neural Polar codes to larger blocklengths and rates.
@@
DeepPolarKO(8,2)Polar(256,37)RM(8,2)Signal-to-noise ratio (SNR) [dB]
@@
Curriculum. We address this by introducing a principled two-stage curriculum training procedure that capitalizes on the inherent nested hierarchy of Polar codes. We leverage two key properties of the Polar coding structure: (1)Hierarchy in k: A Polar$(n,k)$ subsumes all codewords of lower-rate subcodes $\text{Polar}(n,i),1\leq i\leq k$. Curriculum training leveraging the hierarchy in $k$ improved both convergence speed and reliability in the context of neural polar decoding (Hebbar et al., 2023). In the first stage, we use a similar C2N curriculum to pretrain the parameters for each kernel. We progressively train encoder-decoder pairs for DeepPolar ($n=\ell,k=j,\ell)$ codes, for $j=1\to k$ as highlighted in Fig. 4a)), denoted by $g_{\phi}(\ell,j)$, $f_{\theta}(\ell,j)$. (2)Hierarchy in n: In the Plotkin tree, an $\ell\times\ell$ kernel is applied coordinatewise at each level of the Plotkin tree. In the second stage of our curriculum, we initialize each kernel in DeepPolar ($n$,$k$) encoder and decoder with the pre-trained kernels from the first stage, applied at all depths, as illustrated in Fig. 4b). For instance, $g_{12}$ and $g_{20}$, which have three non-frozen inputs, are initialized by a pre-trained kernel $g_{\phi}(4,3)$. Such a systematic and phased training approach significantly enhances the network’s learning efficiency and generalizability (Sec. 4.5).